A charged capacitor and an inductor are connected in series. At time t=0, the current is zero, but the capacitor is fully charged. If T is the period of the resulting oscillations in the LC circuit, the next time after T=0 where the energy stored in the magnetic field of the inductor will be a maximum is:
Answer #1

When t = 0 + T/2, ie after T/2, energy in the inductor is maximum.

The inductor lags behind the voltage by a phase \pi/2, The capacitor leads by \pi/2. So the net phase difference of capacitor with inductor is \pi
A phase difference of \pi = T/2
After a phase difference of \pi, ie T/2, the energy of the inductor will be maximum if the energy of the capacitor were maximum before T/2.